OpenStudy (anonymous):

Choose the best description for the real number square root of 35. A Irrational, because it is not a terminating or repeating decimal Irrational, B because it is a repeating decimal Rational, C because it is not a terminating or repeating decimal D Rational, because it is a repeating decimal

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@annas

OpenStudy (anonymous):

@Valpey

OpenStudy (valpey):

If the square root of 35 were rational, it could be written in the form A/B where A is an integer and B is a whole number. This is how rational numbers are defined.

OpenStudy (anonymous):

Well, the number that you get when square rooting 35 is 5.91608. This wouldn't be a repeating decimal. So, right there you can eliminate the answers that have that.

OpenStudy (valpey):

Also, repeating decimals can all be written as A/B and are therefore rational numbers.

OpenStudy (anonymous):

Then, do like he said, work out if the number is rational or irrational. :)

OpenStudy (valpey):

If it is rational then \(\frac{A}{B}*\frac{A}{B} = 35 = 5*7\) and \(A^2 = 5*7*B^2\). This means that \(A^2\) is a multiple of 5 and a multiple of 7. What does that tell us about A?

OpenStudy (anonymous):

Good job keshaun455! You're correct. The answer is rational. So, which answer would you choose?

OpenStudy (valpey):

??????